OpenStudy (anonymous):

A die is rolled 600 times. The face with six spots appears 112 times. Is the die biased towards that face, or is this just chance variation? The alternative hypothesis is a) The die is biased. b) The chance that the face with six spots appears is greater than 1/6. c) The chance that the face with six spots appears is equal to 112/600. d) The proportion of times the face with six spots appears is equal to 112/600. Answer please! Note: Discard the letter d because i'm sure it is wrong

OpenStudy (amistre64):

we could setup a 99% confidence interval

OpenStudy (anonymous):

generally its 95% C.I. convention

OpenStudy (amistre64):

1/6 \(\pm\) 2.576 sqrt(112*488/600^3)

OpenStudy (amistre64):

1.96 then :)

OpenStudy (amistre64):

or a Pvalue

OpenStudy (anonymous):

1.96 @amistre64

OpenStudy (anonymous):

So the letter is?

OpenStudy (amistre64):

the letter is whatever we get after the process is finished ...

OpenStudy (amistre64):

if you dont understand the process, then ask some questions to learn it

OpenStudy (amistre64):

the Pvalue may be related to: \[\Huge z_{a/2}=\frac{\frac16-\frac{112}{600}}{\sqrt{\frac{112(488)}{600^3}}}\]

OpenStudy (amistre64):

1.26 is within 2 sds

OpenStudy (anonymous):

A and D are definitely wrong..

OpenStudy (amistre64):

it is a wierd question not that i read it again

OpenStudy (amistre64):

*now

OpenStudy (amistre64):

a 95% confidence interval about the sample is from .155 to .218 1/6 = .167 which falls within the range

OpenStudy (anonymous):

Thanks Amistre, Excellent explanation

OpenStudy (amistre64):

youre welcome .... but its still a toss between B and C to me

OpenStudy (amistre64):

1/6 - 112/600 = -.02 so it is greater than 1/6; which is B and the experimental prob is 112/600, which is C

OpenStudy (anonymous):

kropot72 may help... i believe is C but not sure..

OpenStudy (amistre64):

oh ... "the alternative hypothesis is"; i must be blind today

OpenStudy (amistre64):

H1: is equal to 1/6 H2: is not equal to 1/6

OpenStudy (amistre64):

the null has to have an equals in it; the alternative is either >,<, or not=

OpenStudy (amistre64):

this is the only one id choose for an alternative of the options then b) The chance that the face with six spots appears is greater than 1/6.

OpenStudy (anonymous):

Exercise set 5 answers: http://4c51a450.linkbucks.com